You are to roll a fair die n=123 times, each time observing the number of dots appearing on the topside of the die. The number of dots showing on the topside of toss i is a random variable represented by Xi, i=1,2,⋯,123. (a) Consider the distribution of the random variable Xi. Find the mean and the standard deviation of the number of dots showing on the uppermost face of a single roll of this die. μXi= (at least one decimal) σXi= (at least three decimals) (b) Now think about the average of your n=123 tosses of the fair die. What can you say about the distribution of this average? Complete the sentence, using at least three decimals in each numeric answer. The distribution of X with a mean μX= and a standard deviation σX= (c) What is the probability that the average topside on your n=123 tosses will be somewhere between 3.3 and 3.7? Use at least three decimals in your answer. (d) What is the probability that the average of your n=123 tosses is greater than 4? Enter your answer to at least four decimals.
Answer #1

a)

mean =3.5 and the standard deviation =1.7078

b)

distribution of X is approximately normal with a mean μX= 3.5 and a standard deviation σX= 0.1540

c)

probability = P(3.3<X<3.7) = P(-1.299<Z<1.299)= 0.9032-0.0968= 0.8064

(try 0.8060 if this comes wrong)

d)

probability = P(X>4) = P(Z>3.25)= 1-P(Z<3.25)= 1-0.9994= 0.0006